OpenStudy (anonymous):

The lines 3x-7y=-15 and 4x+2y=9 and the x-axis intersect to form a triangle. How many degrees are in the angle of the triangle where the two lines intersect?

OpenStudy (campbell_st):

well is basically the tan formula for the difference start my finding the slope of each line line 1 \[\frac{3}{7}x + \frac{15}{7} = y\] slope m = 3/7 line 2 \[y = -2x + \frac{9}{2}\] slope = -2 then you need \[\tan (\theta) = \frac{m_{1} - m_{2}}{1 + m_{1}m_{2}}\] hope it helps

OpenStudy (campbell_st):

you should also graph the lines... so you can make sense of the answer... as to whether the angle is acute of obtuse